OpenStudy (calculusxy):

When the 69th even natural number is subtracted from the 119th odd natural number, what is the result? @phi @AkashdeepDeb @amistre64 @radar @ranga

OpenStudy (amistre64):

2n is an even number 2n-1 is an odd number assuming natrual is 1,2,3,....

OpenStudy (calculusxy):

Now what do I do?

OpenStudy (amistre64):

well, i would hope that you would define the n=69 even number, and n=119 odd number .. and subtract as needed

OpenStudy (calculusxy):

So is it like 138 (2*69) and 119 (1*119) be like 19?

OpenStudy (amistre64):

2(69), subtracted from 2(119)-1

OpenStudy (calculusxy):

Isn't 2*229-1 greater than 2*69?

OpenStudy (calculusxy):

@amistre64

OpenStudy (ranga):

69th even number is: 69*2 = 138 119th odd number is: (2*119) - 1 = 237 69th even natural number subtracted from the 119th odd natural number is: 237 - 138 = 99.